Calculate the velocity of an electron emitted by radiation of frequency 5.7 x 10^14Hz from a surface with a work function of 3.18 x 10^-19J


1
Expert's answer
2021-08-20T18:04:46-0400

The photoelectric equation is

"\frac{hc}{λ} = ϕ+E_K"

EK=Kinetic energy "=\frac{1}{2}mv^2"

"\frac{hc}{λ} = ϕ+\frac{1}{2}mv^2"

h=plank's constant

c=speed of light

λ=wavelength

ϕ=work function

m=mass of electron

v=velocity of electron

Given

"h=6.626 \times 10^{-34} \;Js \\

c= 3 \times 10^8 \;m/s \\

ϕ = 3.18 \times 10^{-19} \;J \\

v= 5.7 \times 10^{14} \;Hz \\

m= 9.10938 \times 10^{-31} \;kg \\

λ = \frac{c}{frequency} \\

= \frac{3 \times 10^8}{5.7 \times 10^{14}} = 5.26 \times 10^{-7} \;m \\

v = \sqrt{\frac{2}{m} [\frac{hc}{λ }-ϕ] } \\

= \sqrt{\frac{2}{9.10938 \times 10^{-31}} [\frac{6.626 \times 10^{-34} \times 3 \times 10^8}{5.26 \times 10^{-7} }-3.18 \times 10^{-19}] } \\

= \sqrt{2.1955 \times 10^{30} \times (3.779 \times 10^{-19} -3.18 \times 10^{-19} ) } \\

= \sqrt{1.3151 \times 10^{11}} \\

= 3.626 \times 10^5\; m/s"

Answer: "3.626 \times 10^5 \; m/s"